# Learning or Self-aligning? Rethinking Instruction Fine-tuning

### v2

@@ 6 Exp-III: Is Consistency All You Need?
Figure 4 displays the results on three base models under different consistency ratios across four domains. From the figure, we can see that: 1) The performance of the incompatible group (i.e., ratio=0) is indeed poor, which aligns with our previous conclusion; 2) Conversely, relying solely on IFT data that completely aligns with the model’s parameter knowledge (i.e., ratio=1) fails to ensure superior performance across a broad range of scenarios; 3) Optimal performance is most frequently achieved through a balanced integration of incompatible and self-aligning data. The ideal proportion of this combination varies across different foundational models and domains.
@@ 7 Exp-IV: Rethinking Consistency: What Really Matters for IFT?
Figure 3: The regression analysis between the model performance after fine-tuning, and the knowledge consistency between base model and fine-tuned model. We show the results of Mistral-7B in three evaluations.
To this end, we employ a partial correlation assessment Spearman (1961) to analyze the aforementioned Pearson correlation and model performance. Specifically, we utilize a total of 96 models, which include 72 models from Exp-III, in addition to models fine-tuned with consistency ratios of 0.05 and 0.1, conducting tests on homogeneous, in-domain, and out-of-domain data. To eliminate the potential influence of the base model’s performance on the correlation analysis, ensuring that the results only reflect the differences brought about by varying IFT data, we treat the base model’s performance as a control variable. By employing partial correlation assessment under the constraint of removing this control variable’s influence, we analyze the correlation between the Pearson correlation coefficients of model predictions’ rankings before and after IFT and their ultimate performance.
@@
Figure 3 presents the results of grouped regression analysis conducted on Mistral-7b. Table 3 displays the correlation coefficients and significance for the partial correlation analysis across different base models and different evaluations. From the aforementioned figure and table, a significant trend can be observed, namely that the correlation of predictions made by models before and after IFT on a given dataset has a substantial impact on the final performance of the fine-tuned models on that dataset. It is noteworthy that this phenomenon holds true across homogeneous, in-domain, and out-of-domain evaluations. This implies that even for homogeneous and in-domain test sets, which are within the same field as the IFT data, the injection of additional world knowledge and the alteration of internal parameter knowledge through IFT do not contribute to enhancing the final performance of the models. Instead, maintaining consistency in the knowledge of models before and after IFT significantly positively influences the performance of the fine-tuned models.
@@
To further investigate whether this finding is the underlying mechanism responsible for Finding 4, we explore the impact of different consistency data on the correlation of model knowledge before and after IFT in Exp-III. We analyze the probability distribution of model outputs under different IFT data conditions. Table 4 presents the related experimental results, which reveal that: IFT using data containing world knowledge completely inconsistent with the parameter knowledge evidently leads to a divergence in the internal knowledge of the fine-tuned model from that of the original model, thereby impairing the performance of the fine-tuned model. Furthermore, IFT using data containing world knowledge completely consistent with the parameter knowledge also may result in a divergence in the knowledge distribution between the original and fine-tuned models. Specifically, fine-tuning with only consistent IFT data may steer the model towards an ad-hoc knowledge distribution, whereas the original model’s parameter knowledge exists as a relatively smooth distribution. Conversely, using a middle setting that mixes incompatible and self-aligning data allows the optimization process to maintain model parameter knowledge unchanged while preserving the distribution’s smoothness, thereby enabling the fine-tuned model outputs to more closely resemble those of the original model, ultimately yielding better performance.
@@ 8 Conclusion and Discussion
Our discovery not only provides guidance for future IFT data construction, model training, and model evaluation but also provide robust support for some very recent studies. For instance, super-alignment Burns et al. (2023) aims to use a weak model to guide a strong model’s alignment. Our conclusion proves that it is entirely possible to use a weak model with less knowledge to guide a strong, more knowledgeable model for IFT. Our conclusions elucidate the viable underlying factors for such research (Chen et al., 2024;Yuan et al., 2024;Sun et al., 2023;Lu et al., 2024;Zhang et al., 2023;Wan et al., 2024) and provide a solid foundation for the future development of these studies.
@@ A.1 Train
Table 5: The number of instances in the development, testing, and training sets for each domain.
For each item in each train set, we employ five demonstrations from the corresponding test set for in-context learning to assess the parameter knowledge of each base model. The probing prompt is shown in Table 6.
@@
Because model’s responses are influenced by the selection and order of demonstrations in in-context learning (Gao et al., 2023;Min et al., 2022;Wang et al., 2023), we regard responses with a confidence level exceeding 0.5 as reflective of the model’s parameter knowledge to ensure the reliability of identification. For every domain and base model combination, we establish three train datasets of the three distinct settings as described in Section  3. We report the amount of training data for each combination in Table 7. We employ the prompt in Table 8 for constructing instruction-response pairs.
@@
For <explanation> in the prompt, if the answer is consistent with the model’s parameter knowledge but incorrect, we employ prompt in Table 9 for prompting the base model to generate explanations for its choice; if the answer is correct, we directly use the explanation in MedMCQA for medicine and employ prompt in Table 10 for prompting GPT3.5 to generate explanations for the golden answer for the other three domain.
# Instruction Below is a list of conversations between a human and an AI assistant (you). Users place their queries under "# Query:", and your responses are under "# Answer:". You are a helpful, respectful, and honest assistant. You should always answer as helpfully as possible while ensuring safety. Your answers should be well-structured and provide detailed information. They should also have an engaging tone. Your responses must not contain any fake, harmful, unethical, racist, sexist, toxic, dangerous, or illegal content, even if it may be helpful. Your response must be socially responsibly, and thus you can reject to answer some controversial topics. # Query: ```Can you tell me some common types of renewable energy sources? ``` # Answer: ``` Absolutely, below are some of the most common types of renewable energy sources: 1. Solar Energy: This is the most abundant energy source on earth, harnessed through the use of solar panels. These panels convert sunlight into electricity without any moving parts, noise, pollution, or damage to the environment. 2. Wind Energy: Wind turbines convert the kinetic energy in the wind into mechanical power. This mechanical power can be used for specific tasks (such as pumping water) or converted into electricity to power homes, businesses, and schools. 3. Hydropower: Generated by using electricity generators to capture the energy from falling or fast-moving water. This renewable source can come from various ways, including a large dam on a river, tidal or wave energy from the ocean, or using small scale turbines in streams. 4. Geothermal Energy: This type of energy is generated from the heat deep within the Earth. This heat can be used directly for heating buildings or to generate electricity. It is continuously produced inside the Earth and is nearly as reliable as the tides. 5. Biomass Energy: Biomass is organic material that comes from plants and animals, and it contains stored energy from the sun. This energy can be burned directly or converted into biofuel which can burn more efficiently. Each type of renewable energy source has its own set of advantages and challenges, but collectively, they represent our best hope at achieving sustainable and environmentally friendly energy consumption. ``` # Query: ``` Below is a multiple-choice question and the answer. Please give the explanation. Question: <question> Choices: <question options> Answer: <answer> ``` # Answer:
@@
For stable and fair IFT, we incorporate an equal proportion of general instruction data sampled from alpaca-gpt4-en (Peng et al., 2023) which is identical for each domain and base model combination. We use vicuna-v1.5 Zheng et al. (2023) format to train models.
@@ A.2 Test
To compare the performance across three settings for every domain and base model combination, we devise three evaluation types: homogeneous, in-domain, and out-of-domain tests. The first evaluation involves a holdout from the same multi-choice question set, while the subsequent two, in-domain and out-of-domain, are derived from MMLU’s splits. The specific details of the splits for each domain are detailed in Table 11.
@@
Engineering
["electrical_engineering"]
History
["high_school_european_history", "high_school_us_history", "high_school_world_history", "prehistory"]
Jurisprudence
["econometrics", "high_school_geography", "high_school_government_and_politics", "high_school_macroeconomics", "high_school_microeconomics", "high_school_psychology", "human_sexuality", "international_law", "jurisprudence", "professional_law", "sociology", "public_relations", "professional_psychology", "security_studies", "us_foreign_policy"]
Medicine
["anatomy", "clinical_knowledge", "college_medicine", "human_aging", "medical_genetics", "nutrition", "professional_medicine", "virology"]

### v3

@@ 6 Exp-III: Is Consistency All You Need?
Figure 4 displays the results on three base models under different consistency ratios across four domains. From the figure, we can see that: 1) The performance of the incompatible group (i.e., ratio=0) is indeed poor, which aligns with our previous conclusion; 2) Conversely, relying solely on IFT data that completely aligns with the model’s parameter knowledge (i.e., ratio=1) fails to ensure superior performance across a broad range of scenarios; 3) Optimal performance is most frequently achieved through a balanced integration of incompatible and self-aligning data. The ideal proportion of this combination varies across different base models and domains.
@@ 7 Exp-IV: Rethinking Consistency: What Really Matters for IFT?
Figure 3: The regression analysis between the model performance after fine-tuning, and the knowledge consistency between base model and fine-tuned model. We show the results of Mistral-7B in three evaluations. The grouped linear regression demonstrate the positive correlations between the model performance after IFT and model internal knowledge consistency before and after IFT. Points in the same regression line indicate the results of the same base model fine-tuned with different IFT data of the same domain on the same test set (HOMO, ID, or OOD).
To this end, we employ a partial correlation assessment Spearman (1961) to analyze the aforementioned Pearson correlation and model performance. Specifically, we utilize a total of 96 models, which include 72 models from Exp-III, in addition to models fine-tuned with consistency ratios of 0.05 and 0.1, conducting tests on homogeneous, in-domain, and out-of-domain data. To eliminate the potential influence of the base model’s performance on the correlation analysis, ensuring that the results only reflect the differences brought about by varying IFT data, we treat the base model’s performance as a control variable. By employing partial correlation assessment under the constraint of removing this control variable’s influence, we analyze the correlation between the Pearson correlation coefficients of model predictions’ rankings before and after IFT and their ultimate performance. Please refer to the Appendix D for more details about our analysis.
@@
Figure 3 presents the results of grouped regression analysis conducted on Mistral-7b. Table 3 displays the correlation coefficients and significance for the partial correlation analysis across different base models and different evaluations. From the aforementioned figure and table, a significant trend can be observed, namely that the correlation of predictions made by models before and after IFT on a given evaluation has a substantial impact on the final performance of the fine-tuned models on that evaluation. It is noteworthy that this phenomenon holds true across homogeneous, in-domain, and out-of-domain evaluations. This implies that even for homogeneous and in-domain test sets, which are within the same field as the IFT data, the injection of additional world knowledge and the alteration of internal parameter knowledge through IFT do not contribute to enhancing the final performance of the models. Instead, maintaining consistency in the knowledge of models before and after IFT significantly positively influences the performance of the fine-tuned models.
@@
To further investigate whether this finding is the underlying mechanism responsible for Finding 4, we explore the impact of different consistency data on the correlation of model knowledge before and after IFT in Exp-III. We analyze the probability distribution of model outputs under different IFT data conditions. Table 4 presents the related experimental results, which reveal that: IFT using data containing world knowledge completely inconsistent with the parameter knowledge evidently leads to a divergence in the internal knowledge of the fine-tuned model from that of the original model, thereby impairing the performance of the fine-tuned model. Furthermore, IFT using data containing world knowledge completely consistent with the parameter knowledge also may result in a divergence in the knowledge distribution between the original and fine-tuned models. Specifically, fine-tuning with only consistent IFT data may steer the model towards an sharp knowledge distribution, whereas the original model’s parameter knowledge exists as a relatively smooth distribution. Conversely, using a middle setting that mixes incompatible and self-aligning data allows the optimization process to maintain model parameter knowledge unchanged while preserving the distribution’s smoothness, thereby enabling the fine-tuned model outputs to more closely resemble those of the original model, ultimately yielding better performance.
@@ 8 Conclusion and Discussion
Our discovery not only provides guidance for future IFT data construction, model training, and model evaluation but also provide robust support for some very recent studies. For instance, super-alignment Burns et al. (2023) aims to use a weak model to guide a strong model’s alignment. Our conclusion proves that it is entirely possible to use a weak model with less knowledge to guide a strong, more knowledgeable model for IFT. Our conclusions elucidate the viable underlying factors for self-instruction-tuning Sun et al. (2024); Guo et al. (2024), self-rewarding Yuan et al. (2024) and consistent alignment Wan et al. (2024) and provide a solid foundation for the future development of these studies.
@@
Acknowledgements
We sincerely thank all anonymous reviewers for their insightful comments and valuable suggestions. This research work is supported by CAS Project for Young Scientists in Basic Research (Grant No.YSBR-040), the National Natural Science Foundation of China under Grants no. 62122077 and 62106251, and the Basic Research Program of ISCAS, Grant No. ISCAS-JCZD-202303.
@@ A.1 Train
Table 5: The number of instances in the development, test, and train sets for each domain.
For each item in each train set, we employ five demonstrations from the corresponding dev set for in-context learning to assess the parameter knowledge of each base model. The probing prompt is shown in Table 6.
@@
Because model’s responses are influenced by the selection and order of demonstrations in in-context learning Gao et al. (2023); Min et al. (2022), we regard responses with a confidence level exceeding 0.5 as reflective of the model’s parameter knowledge to ensure the reliability of identification. For every domain and base model combination, we establish train datasets on the three different settings as described in Section  3. We report the amount of training data for each combination in Table 7. We employ the prompt in Table 8 for constructing instruction-response pairs.
@@
For <explanation> in the prompt, if the answer is consistent with the model’s parameter knowledge but incorrect, we employ prompt in Table 9 for prompting the base model to generate explanations for its choice; if the answer is correct, we directly use the explanation in MedMCQA for medicine and employ prompt in Table 10 for prompting GPT3.5 to generate explanations for the golden answer for the other three domains.
# Instruction Below is a list of conversations between a human and an AI assistant (you). Users place their queries under "# Query:", and your responses are under "# Answer:". You are a helpful, respectful, and honest assistant. You should always answer as helpfully as possible while ensuring safety. Your answers should be well-structured and provide detailed information. They should also have an engaging tone. Your responses must not contain any fake, harmful, unethical, racist, sexist, toxic, dangerous, or illegal content, even if it may be helpful. Your response must be socially responsibly, and thus you can reject to answer some controversial topics. # Query:
```Can you tell me some common types of renewable energy sources? ```
# Answer:
Absolutely, below are some of the most common types of renewable energy sources: 1. Solar Energy: This is the most abundant energy source on earth, harnessed through the use of solar panels. These panels convert sunlight into electricity without any moving parts, noise, pollution, or damage to the environment. 2. Wind Energy: Wind turbines convert the kinetic energy in the wind into mechanical power. This mechanical power can be used for specific tasks (such as pumping water) or converted into electricity to power homes, businesses, and schools. 3. Hydropower: Generated by using electricity generators to capture the energy from falling or fast-moving water. This renewable source can come from various ways, including a large dam on a river, tidal or wave energy from the ocean, or using small scale turbines in streams. 4. Geothermal Energy: This type of energy is generated from the heat deep within the Earth. This heat can be used directly for heating buildings or to generate electricity. It is continuously produced inside the Earth and is nearly as reliable as the tides. 5. Biomass Energy: Biomass is organic material that comes from plants and animals, and it contains stored energy from the sun. This energy can be burned directly or converted into biofuel which can burn more efficiently. Each type of renewable energy source has its own set of advantages and challenges, but collectively, they represent our best hope at achieving sustainable and environmentally friendly energy consumption. ``` # Query: ``` Below is a multiple-choice question and the answer. Please give the explanation. Question: <question> Choices: <question options> Answer: <answer> ``` # Answer:
@@
For stable and real IFT, we incorporate an equal proportion of general instruction data sampled from alpaca-gpt4-en Peng et al. (2023) which is identical for each domain and base model combination. We use vicuna-v1.5 Zheng et al. (2024) format to train models.
@@ A.2 Test
To compare the performance across three settings for each domain and base model combination, we devise three evaluation types: homogeneous, in-domain, and out-of-domain tests. The first evaluation involves a holdout from the same multi-choice question set, while the subsequent two, in-domain and out-of-domain, are derived from MMLU’s splits. The specific details of the splits for each domain are detailed in Table 11.
@@
Engineering ["electrical_engineering"]
History ["high_school_european_history", "high_school_us_history", "high_school_world_history", "prehistory"]
Jurisprudence ["econometrics", "high_school_geography", "high_school_government_and_politics", "high_school_macroeconomics", "high_school_microeconomics", "high_school_psychology", "human_sexuality", "international_law", "jurisprudence", "professional_law", "sociology", "public_relations", "professional_psychology", "security_studies", "us_foreign_policy"]
Medicine ["anatomy", "clinical_knowledge", "college_medicine", "human_aging", "medical_genetics", "nutrition", "professional_medicine", "virology"]
@@ Appendix C Performance of Models Fine-tuned in Different Consistency Ratios
Appendix D Details about Spearman Partial Correlation Analysis
Partial correlation measures the degree of association between two variables (here, the model performance after IFT and the knowledge consistency between the base model and fine-tuned model) while controlling the effect of one or more additional variables (the accuracy on the same test set of the base model in our case).
To calculate the Spearman partial correlation between the model performance after IFT and the knowledge consistency between the base model and fine-tuned model, we first calculate the knowledge consistency, measured by the correlation coefficient of model prediction ranking before and after IFT using the following formula.
$pr=\frac{1}{n}\sum_{i=1}^{|D_{\text{test}}|}\text{pearson}(\text{sort}(m_{\text{base}},q_{i},c_{i}),\text{sort}(m_{\text{tuned}},q_{i},c_{i}))$
where $q_{i}$ and $c_{i}$ represent the i-th question and its corresponding set of candidate choices. The terms $m_{\text{base}}$ and $m_{\text{tuned}}$ denote the base model and its fine-tuned model, respectively. $D_{\text{test}}$ represents any test set like HOMO, ID, or OOD. $|D_{\text{test}}|$ denotes the size of the test set. The ‘sort‘ function ranks the candidate choices based on the predicted probability distribution from the input model. $pr$ reflects the average Pearson correlation coefficient between the rankings predicted by the base model and the fine-tuned model of each question.
For each base model, its every fine-tuned model, and each test set, we calculate the corresponding above $pr$ and organize them along with the corresponding accuracy on the same test set of the base model and fine-tuned model to three lists, denoted as $pr_{s}$, $p_{\text{base}}$ and $p_{\text{tuned}}$. Then we compute the Spearman partial correlation. The formulas are:
$e(pr_{s})$ $=pr_{s}-\hat{pr_{s}}(p_{\text{base}})$
$e(p_{\text{tuned}})$ $=p_{\text{tuned}}-\hat{p_{\text{tuned}}}(p_{\text{base}})$
$r,p\text{-}value$ $=\text{spearman}(e(pr_{s}),e(p_{\text{tuned}}))$
Here, $\hat{pr_{s}}(p_{\text{base}})$ and $\hat{p_{\text{tuned}}}(p_{\text{base}})$ present the estimated value of $pr_{s}$ and $p_{\text{tuned}}$ predicted based on $p_{\text{base}}$, respectively. The residuals of $pr_{s}$ and $p_{\text{tuned}}$ relative to $p_{\text{base}}$ are denoted as $e(pr_{s})$ and $e(p_{\text{tuned}})$, respectively. We finally calculate Spearman’s correlation coefficient and the corresponding p-values between the residuals, denoted as $r$ and $p\text{-}value$.
In this way, we can find the direct relationship between the model performance after IFT, and the knowledge consistency between the base model and fine-tuned model by removing the influence of base model performance that affects them both.